# SALSA-RL: Stability Analysis in the Latent Space of Actions for Reinforcement Learning

### v1

@@ Abstract
Modern deep reinforcement learning (DRL) methods have made significant advances in handling continuous action spaces. However, real-world control systems–especially those requiring precise and reliable performance–often demand formal stability, and existing DRL approaches typically lack explicit mechanisms to ensure or analyze stability. To address this limitation, we propose SALSA-RL (Stability Analysis in the Latent Space of Actions), a novel RL framework that models control actions as dynamic, time-dependent variables evolving within a latent space. By employing a pre-trained encoder-decoder and a state-dependent linear system, our approach enables both stability analysis and interpretability. We demonstrated that SALSA-RL can be deployed in a non-invasive manner for assessing the local stability of actions from pretrained RL agents without compromising on performance across diverse benchmark environments. By enabling a more interpretable analysis of action generation, SALSA-RL provides a powerful tool for advancing the design, analysis, and theoretical understanding of RL systems.
Keywords Reinforcement learning  $\cdot$ Dynamical system  $\cdot$ Stability analysis
@@ 1 Introduction
Figure 1: Overview of the SALSA-RL framework. Our proposed augmentation to RL algorithms relies on a latent action representation governed by a time-varying linear dynamical system governed by a state-conditioned matrix $\mathbf{A}_{t}$. This enables local stability analyses in the action-state phase for reliable and interpretable RL deployments. The framework integrates seamlessly with existing DRL algorithms, maintaining competitive performance. The contour shows how latent action stability evolves, with the policy driving the system from instability to stability. Stability boundaries (white) mark critical state regions where large deviations of angular shifts, can cause instability or loss of control.
@@
Additionally, observing how actions evolve provides insight into a policy’s ability to maintain stability by avoiding erratic decision patterns, adapting to disturbances with smooth control adjustments, and balancing exploration and exploitation for effective regulation without excessive correction. This analysis also enhances RL interpretability through various different aspects:
@@
    Stability: Are control actions and observations (states) able to achieve stability within local or global regions?
@@
    Predictability and Consistency: Do action patterns follow a logical and predictable progression, across similar conditions?
  3. 3.
    Risk Identification: Can unsafe or failure-prone state-action combinations be identified through the interaction?
Identifying such patterns not only aids in policy validation but also helps detect subtle issues like overcompensation, delayed reactions, or unstable oscillations that may not be evident when observing state trajectories alone.
In light of the importance of action dynamics for understanding control strategies, we propose SALSA-RL, a novel RL framework for dynamical system control that leverages action dynamics to analyze and design stable control strategies while enhancing interpretability. Unlike standard RL approaches where actions are treated as discrete decisions, our framework models actions in a latent space, evolving continuously based on the state information. We achieve this by employing a pre-trained autoencoder for action encoding and decoding. The action representation is encoded into a latent space, where the latent action dynamics $z_{t}$ evolve following a differential equation of the form: $\dot{z}=\mathbf{A}_{t}$, where $\mathbf{A}_{t}$ is a state-dependent dynamic matrix, learned through a deep neural network conditioned on the current state. This structured representation not only allows for local-stability analysis of control policies but also provides a means to visualize and interpret the constraints imposed on the state space, enhancing both policy transparency and generalization. Figure 1 illustrates the details of SALSA-RL.
This proposed framework is compatible with existing popular DRL methods (such as SAC, DDPG, TD3 [14, 11, 27]) and can be easily integrated using exactly the same standard training strategies while maintaining competitive control performance. Beyond achieving effective control, SALSA-RL provides valuable tools for evaluating trained policies. Since actions directly influence state evolution, the influence of the action on local stability of action dynamics serves as an indirect yet effective indicator of the stability of the state trajectories. Furthermore, as a key variable in the latent dynamics, the time-dependent square matrix $\mathbf{A}_{t}$, allows for the incorporation of local-in-time eigenvalue-based stability analysis. Consequently, one can perform a rigorous evaluation of the system’s behavior. This analysis can provide insights into how action-space and constraints impact long-term stability and how state transitions behave near critical regions, further enhancing the interpretability of the learned control policy.
@@
    We propose SALSA-RL, a framework that models control actions as continuous dynamics evolving in the latent space with a time-varying linear system, enabling a deeper understanding of control strategies and dynamics.
@@
    Our stability analysis demonstrates how the control policy achieves stability in states and actions while identifying and avoiding high-risk (unstable) regions.
  4. 4.
    We show how key states affect system stability and control actions, potentially providing insights to inform better reward design and policy refinement.
@@ 2.2 Interpretability in DRL
2.3 RL for Dynamical System Controls
DRL is well-studied in physical applications, such as optimizing flow control and turbulent fluid dynamics, where domain knowledge is often available, and state-control actions exhibit stronger correlations compared to other domains[47, 43, 12, 34]. Recent approaches have further improved interpretability by leveraging state-action correlations and domain knowledge.
Symbolic and Neural-Guided Approaches Neural-guided methods like NUDGE [8], PIRL [41], and NLRL [18] use symbolic reasoning to discover interpretable policies [19]. Symbolic controllers [23, 36], such as DSP [26] and SINDy-RL [50], employ explicit state-control mappings to derive generalizable control laws. These methods improve interpretability and generalizability, making them suitable for safe, explainable decision-making. However, neural-guided approaches rely on logical representations [8], limiting temporal precision and scalability in complex systems. Similarly, symbolic controllers face scalability issues in multi-dimensional or stochastic tasks where predefined forms fail to capture intricate interactions [26]. While explicit formulations aid stability analysis, they may overlook subtle instabilities in evolving or poorly understood dynamics.
@@ 3.1 Problem Formulation
Assuming the Markov property, the next state $s_{t+1}$ depends only on the current state $s_{t}$ and action $a_{t}$, governed by the transition probability $P(s_{t+1}\mid s_{t},a_{t})$. The agent maximizes the cumulative discounted reward $R=\sum_{t=0}^{T-1}\gamma^{t}r(s_{t},a_{t})$, where $\gamma\in[0,1]$ balances immediate and future rewards over the time horizon $T$.
@@
where $\tau=(\mathbf{s}_{0},\mathbf{a}_{0},\dots,\mathbf{s}_{T},\mathbf{a}_{T})$ represents trajectories sampled from the policy. Although the latent action evolves in continuous time, the discrete-time MDP formulation is sufficient for the proposed algorithm, as actions are applied at discrete intervals $\Delta t$.
@@ 3.2 SALSA-RL Control Framework
The proposed framework aims to simplify time-dependent dynamical system control by encoding actions into a compact latent space, where state-dependent transformations govern their evolution. The framework consists of the following:
@@
where $\mathbf{z}_{t}\in\mathbb{R}^{d_{h}}$ is the latent representation of the action, $d_{h}$ is the latent dimension size, and $\mathbf{a}_{0}=\mathbf{0}$.
@@
where $\mathbf{A}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ is a state-dependent matrix learned by the neural network $\mathcal{N}_{\theta}$ conditioned on the current state.
@@
  Input: Initial state $\mathbf{s}_{0}$, initial action $\mathbf{a}_{0}=\mathbf{0}$, time horizon $T$, pre-trained Encoder and Decoder
  $\mathbf{z}_{0}\leftarrow\textbf{Encoder}(\mathbf{a}_{0})$ // Encode action
@@
     $\mathbf{a}_{t+1}\leftarrow\textbf{Decoder}(\mathbf{z}_{t+1})$ // Decode latent to next action
     $\mathbf{s}_{t+1},Done\leftarrow\text{env.step}(\mathbf{a}_{t+1})$
     $\mathbf{a}_{t}\leftarrow\mathbf{a}_{t+1}$
     if $Done$ then
@@ 3.3 Training Procedure
The training process consists of two stages: (1) pretraining the encoder-decoder network and (2) optimizing the latent dynamic module.
Encoder-Decoder Network. This network is trained to create a compact and effective latent representation of actions, optimized by minimizing the reconstruction loss:
@@
where $\hat{\mathbf{a}}_{t}$ is the reconstructed action from the decoder. To train the module, training data, actions, are generated using trained DRL policies such as SAC and TD3. This data is then used to train the action encoding and decoding modules. We detailed this procedure in Appendix C.
Latent Dynamic Policy. Following this, the proposed policy is initialized in a similar manner as most DRL methods for continuous action control, such as PPO [37], SAC [14], DDPG [27], and TD3 [11]. The policy network (actor), typically generating actions in DRL as $\mathbf{a}_{t}=\mathcal{N}_{\theta}(\mathbf{s}_{t})$, is instead employed to predict a state-dependent dynamic matrix $\mathbf{A}_{t}=\mathcal{N}_{\theta}(\mathbf{s}_{t},\mathbf{a}_{t-1})$. The gradients of the objective to maximize the expected cumulative reward, in deterministic policies, for example, can be expressed generally as:
$\nabla_{\theta}J(\pi_{\theta})\;=\;\mathbb{E}_{\tau\sim\pi_{\theta}(\mathbf{a}_{t}\mid\mathbf{s}_{t},\mathbf{a}_{t-1})}\left[\sum_{t=0}^{T-1}\gamma^{t}\nabla_{\theta}R\bigl{(}\mathbf{s}_{t},\mathbf{a}_{t}\bigr{)}\right]$ (8)
Meanwhile, the critic network structure, if available, remains unchanged, retaining its role in evaluating the state-action value function. This design allows SALSA-RL to leverage existing DRL algorithms while providing flexibility to represent and optimize latent dynamics. A detailed gradient computation is provided in Appendix A.
@@ 3.4 Stability Analysis and Interpretability
Eigenvalue-based Stability. To assess the stability of learned control policies, we apply eigenvalue-based analysis to the dynamic matrix $\mathbf{A}_{t}$. Stability at each time step is determined by the magnitude of its largest eigenvalue $\lambda_{1}$.
$\lambda_{1}=\max\left(\text{eigvals}(\mathbf{A}_{t})\right).$ (9)
A policy is considered locally stable if the largest eigenvalue satisfies $\mathrm{Re}(\lambda_{1})<1$, ensuring that latent dynamics do not diverge over time. Additionally, a non-zero imaginary component $\mathrm{Im}(\lambda_{1})$ indicates oscillatory behavior. Specifically: $\mathrm{Re}(\lambda_{1})<1$ and $\mathrm{Im}(\lambda_{1})\neq 0$ indicates damped oscillations, $\mathrm{Re}(\lambda_{1})=0$ and $\mathrm{Im}(\lambda_{1})\neq 0$ indicates pure oscillations, and $\mathrm{Re}(\lambda_{1})>1$ implies unstable dynamics.
Additionally, the neighborhood states around the observed states, denoted as $s_{\text{range}}$, are evaluated by computing the corresponding $\mathbf{A}_{t}$ and analyzing the eigenvalues. This evaluation offers an overview of stable and unstable regions during control, capturing local state landscapes that may influence stability. It provides insights into regions where the policy achieves stable control or displays unstable behavior. The procedure is outlined in Algorithm 2.
Algorithm 2 Latent Dynamic Eigenvalue Evaluation
  Input: Observed $n$ states $s_{t}^{(i)}$, where $i=1,2,...,n$
  for $i=1$ to $n$ do
     for $t=0$ to $T$ do
        for $s$ in $s_{\text{range}}^{(i)}$ do
           $\hat{\mathbf{s}}[i]\leftarrow s$ // Replace with neighboring values
           $\mathbf{A}_{t}\leftarrow\mathcal{N}_{\theta}(\hat{\mathbf{s}})$
           $\lambda_{1}\leftarrow\text{max}\bigl{(}\text{eigvals}(\mathbf{A}_{t})\bigl{)}$ // First eigenvalue
        end for
     end for
  end for

### v3

@@ Abstract
Modern deep reinforcement learning (DRL) methods have made significant advances in handling continuous action spaces. However, real-world control systems–especially those requiring precise and reliable performance–often demand interpretability in the sense of a-priori assessments of agent behavior to identify safe or failure-prone interactions with environments. To address this limitation, we propose SALSA-RL (Stability Analysis in the Latent Space of Actions), a novel RL framework that models control actions as dynamic, time-dependent variables evolving within a latent space. By employing a pre-trained encoder-decoder and a state-dependent linear system, our approach enables interpretability through local stability analysis, where instantaneous growth in action-norms can be predicted before their execution. We demonstrate that SALSA-RL can be deployed in a non-invasive manner for assessing the local stability of actions from pretrained RL agents without compromising on performance across diverse benchmark environments. By enabling a more interpretable analysis of action generation, SALSA-RL provides a powerful tool for advancing the design, analysis, and theoretical understanding of RL systems.
@@ 1 Introduction
Figure 1: Overview of the SALSA-RL framework. Our proposed augmentation to pre-trained RL algorithms relies on a latent action representation governed by a time-varying linear dynamical system through a state-conditioned matrix $\mathbf{A}_{t}$. This enables local stability analyses in the action-state phase for reliable and interpretable RL deployments. The framework integrates seamlessly with existing RL algorithms, maintaining competitive performance. The contours represent the dynamically changing spectral radius of the latent linear system, with the policy seeking bounded regions (in white) with high local stability. Consequently, initializing controllers in regions outside these regions leads to high-risk behavior and potential failure.
@@
Additionally, observing how actions evolve provides insight into a policy’s ability to maintain stability by avoiding erratic decision patterns, adapting to disturbances with smooth control adjustments, and balancing exploration and exploitation for effective regulation without excessive correction. This analysis also enhances RL interpretability through an answer to the following questions:
@@
    Given pretrained reinforcement learning agents, can we identify regions of smooth or regular agent behavior in action-state phase space?
@@
    Conversely, can we also identify regions with a high degree of discontinuous (and consequently high-risk/failure-prone) behavior?
Identifying such patterns not only aids in policy validation but also helps detect subtle issues like overcompensation, delayed reactions, or unstable oscillations that may not be evident when observing state or action trajectories in their original high-dimensional spaces. Moreover, in many physical systems, actions are often the only fully observable or controllable signals, whereas the state may be difficult or impractical to measure directly. In addition, actions are typically subject to physical constraints or actuation budgets that may not be fully embedded during learning. Anticipating how actions behave, especially under changing conditions, can provide valuable insights for safety and planning.
In light of the importance of action dynamics for understanding control strategies, we propose SALSA-RL, a novel RL framework for dynamical system control that leverages action dynamics to analyze and design stable control strategies while enhancing interpretability. Unlike standard RL approaches, where actions are treated as discrete decisions, our framework models actions in a latent space, evolving based on the state information. We achieve this by employing a pre-trained autoencoder for action encoding and decoding. The action representation is encoded into a latent space, where the latent action dynamics $z_{t}$ evolve following a differential equation of the form: $\dot{z}=\mathbf{A}_{t}z_{t}$, where $\mathbf{A}_{t}$ is a state-dependent dynamic matrix, learned through a deep neural network conditioned on the current state. This allows for local stability analysis of action dynamics in the latent space. In particular, we define local stability in the context of a state-dependent and time-varying linear system. Stability analysis of such a system provides information about the short-term growth of action-dynamic norms. Moreover, our structured representation not only allows for local-stability analysis of control policies but also provides a means to visualize and interpret the constraints imposed on the state space, enhancing both policy transparency and generalization. Figure 1 illustrates the details of SALSA-RL.
This proposed framework is compatible with existing popular DRL methods (such as SAC, DDPG, TD3 [16, 12, 32]) and can be easily integrated using exactly the same standard training strategies while maintaining competitive control performance. Beyond achieving effective control, SALSA-RL provides valuable tools for evaluating trained policies. Since actions directly influence state evolution, the influence of the action on the local stability of action dynamics serves as an indirect yet effective indicator of the stability of the state trajectories. Furthermore, as a key variable in the latent dynamics, the time-dependent square matrix $\mathbf{A}_{t}$ allows for the incorporation of local-in-time eigenvalue-based stability analysis. Consequently, one can perform a rigorous evaluation of the system’s behavior. This analysis can provide insights into how action-space and constraints impact long-term stability and how state transitions behave near critical regions, further enhancing the interpretability of the learned control policy.
@@
    We propose SALSA-RL, a framework that models control actions as discrete-time dynamics in latent space via a time-varying linear system, enabling deeper insight into control strategies and dynamics.
@@
    We perform local stability analysis to identify regions of action-state phase space for regular and high-risk behavior. We also demonstrate that pretrained agents seek out regions of regularity.
@@ 2.2 Interpretability in DRL
2.3 Dynamical System Control and Deep Reinforcement Learning
Methods such as Koopman-based control [35, 57] and Embed-to-Control [51] model nonlinear dynamics through latent linearization, where the system is approximated as linear in a lifted latent space. Though not inherently RL, these methods have inspired RL frameworks with structured latent dynamics that often rely on globally fixed linear operators—rather than state-dependent ones—thereby limiting adaptability to state-dependent variations in very high-dimensional nonlinear dynamical systems [53, 44]. In particular, [44] introduces value function learning with dictionary-based methods that can provide increased insight into agent behavior. However, a direct connection to the inherent local regularity of the trained agent’s behavior is not clear.
On the other hand, DRL is well-studied in physical applications, such as optimizing flow control and turbulent fluid dynamics, where domain knowledge is often available, and state-control actions exhibit stronger correlations compared to other domains[58, 52, 13, 41]. Recent approaches (summarized below) have further improved interpretability by leveraging state-action correlations and domain knowledge.
Symbolic and Neural-Guided Approaches Neural-guided methods like NUDGE [9], PIRL [49], and NLRL [21] use symbolic reasoning to discover interpretable policies [23]. Symbolic controllers [28, 43], such as DSP [31] and SINDy-RL [62], employ explicit state-control mappings to derive generalizable control laws. These methods improve interpretability and generalizability, making them suitable for safe, explainable decision-making. However, neural-guided approaches rely on logical representations [9], limiting temporal precision and scalability in complex systems. Similarly, symbolic controllers face scalability issues in multi-dimensional or stochastic tasks where predefined forms fail to capture intricate interactions [31]. While explicit formulations may aid analysis of agent behavior, explicit stability assessments for complex high-dimensional learning tasks are challenging.
@@ 3.1 Problem Formulation
Assuming the Markov property, the next state $s_{t+1}$ depends on the current state $s_{t}$ and action $a_{t}$, governed by the transition probability $P(s_{t+1}\mid s_{t},a_{t})$. The agent maximizes the cumulative discounted reward $R=\sum_{t=0}^{T-1}\gamma^{t}r(s_{t},a_{t})$, where $\gamma\in[0,1]$ balances immediate and future rewards over the time horizon $T$.
@@
where $\tau=(\mathbf{s}_{0},\mathbf{a}_{0},\dots,\mathbf{s}_{T},\mathbf{a}_{T})$ represents trajectories sampled from the policy. Although the latent action $z$ evolves in continuous time, the discrete-time MDP formulation is sufficient for the proposed algorithm, as actions are applied at discrete intervals $\Delta t$.
@@ 3.2 SALSA-RL Control Framework
The proposed framework simplifies control of time-dependent dynamical systems by encoding actions into a compact latent space, where state-dependent transformations govern their evolution. It consists of:
@@
where $\mathbf{z}_{t}\in\mathbb{R}^{d_{h}}$ is the latent representation of the action, $d_{h}$ is the latent dimension size, and $\mathbf{a}_{0}=\mathbf{0}$. Here, we intentionally project actions into a higher-dimensional latent space (typically $h_{d}>\dim(a)$) to provide greater expressive capacity for learning structured dynamics. This higher-dimensional representation eases training and supports more interpretable stability analysis, especially in complex or underactuated tasks.
@@
where $\mathbf{A}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ is a state-dependent matrix learned by the neural network $\mathcal{N}_{\theta}$ conditioned on the current state. Importantly, to ensure numerical stability, we apply a $\tanh$ activation function followed by a scaling factor of 2 to the output of the latent dynamics network, thereby constraining all elements of $A_{t}$ within the range [2,2][-2,2]. While we also experimented with diagonal forms of $A_{t}$ to simplify the analysis, we found that full (dense) matrices enabled richer inter-dependencies and better performance in control.
It is worth noting that we introduce a latent action space not for dimensionality reduction but to enable structured analysis of action behavior in the state–action phase space. This design empirically yields smoother dynamics and more stable training than operating directly in the original action space, while also allowing flexible dimensionality for constructing square dynamics matrices. It further aligns with Koopman operator theory, where systems are lifted to higher-dimensional spaces for structured modeling before being mapped back.
@@
In practice, we decode $\mathbf{z}_{t+1}$ into an action $a_{t+1}$ for system control. Then, the same $a_{t+1}$ is encoded back into a new latent vector $\mathbf{z}_{t+1}^{\prime}$. This mimics real-world control scenarios where only the executed action is observable, and the internal latent representation must be reconstructed based on the current state and action. This stepwise decode-encode process also reflects how a deployed agent operates under partial observability or noisy dynamics, maintaining consistency through latent transitions. Additionally, in line with standard RL environment setups, the decoded action is clipped in all our experiments. Importantly, this clipping is applied only after decoding; the latent space itself remains unconstrained.
@@
Input: Initial state $\mathbf{s}_{0}$, initial action $\mathbf{a}_{-1}=\mathbf{0}$, time horizon $T$, pre-trained Encoder and Decoder
@@
  $\mathbf{z}_{t}\leftarrow\textbf{Encoder}(\mathbf{a}_{t-1})$ // Encode action
@@
  $\mathbf{a}_{t}\leftarrow\textbf{Decoder}(\mathbf{z}_{t+1})$ // Decode latent to next action
  $\mathbf{s}_{t+1},Done\leftarrow\text{env.step}(\mathbf{a}_{t})$
  $\mathbf{a}_{t-1}\leftarrow\mathbf{a}_{t}$
  if DoneDone then
@@ 3.3 Training Procedure
Training involves two stages: (1) encoder-decoder pretraining, and (2) latent dynamics optimization. Encoder-Decoder Network. This network is trained to create a compact and effective latent representation of actions, optimized by minimizing the reconstruction loss:
@@
where $\hat{\mathbf{a}}_{t}$ is the reconstructed action from the decoder. To train the module, training data $D$, actions, are constructed from actions drawn either uniformly over the full action space or collected from pretrained agent. This data is then used to train the action encoding and decoding modules. We detailed this procedure in Appendix C.
Latent Dynamic Policy. Following this, the proposed policy is initialized in a similar manner as most DRL methods for continuous action control, such as PPO [45], SAC, DDPG, and TD3. The policy network (actor), typically generating actions in DRL as $\mathbf{a}_{t}=\mathcal{N}_{\theta}(\mathbf{s}_{t})$, is instead employed to predict a state-dependent dynamic matrix $\mathbf{A}_{t}=\mathcal{N}_{\theta}(\mathbf{s}_{t},\mathbf{a}_{t-1})$. The gradients of the objective to maximize the expected cumulative reward, in deterministic policies, for example, can be expressed generally as:
$\hskip-2.84526pt\nabla_{\theta}J(\pi_{\theta})\;=\;\mathbb{E}_{\tau\sim\pi_{\theta}(\mathbf{a}_{t}\mid\mathbf{s}_{t},\mathbf{a}_{t-1})}\left[\sum_{t=0}^{T-1}\gamma^{t}\nabla_{\theta}R\bigl{(}\mathbf{s}_{t},\mathbf{a}_{t}\bigr{)}\right]$ (8)
Meanwhile, the critic network architecture, if available, remains unchanged, evaluating the state-action value function. This design allows SALSA-RL to leverage existing DRL algorithms while providing flexibility to represent and optimize latent dynamics. A detailed gradient computation is provided in Appendix A.
@@ 3.4 Stability Analysis and Interpretability
Local Stability Analysis. While the environments are continuous-time, the action dynamics modeled in SALSA-RL follow discrete-time updates. Therefore, local stability is assessed using the spectral radius of the state-dependent linear term that evolves agent actions in the latent space. The stability analysis module in SALSA-RL is applied after training and does not interfere with the training process itself. Thus, it serves as a post-hoc diagnostic tool for pretrained policies, rather than modifying the optimization pipeline. At each time step, we compute:
$\rho(\mathbf{A}_{t})=\max_{i}\left|\lambda_{i}(\mathbf{A}_{t})\right|,$ (9)
where $\lambda_{i}(\mathbf{A}_{t})$ are the eigenvalues of the state-dependent matrix $\mathbf{A}_{t}$. Here, local stability is guaranteed when all eigenvalues of $\mathbf{A}_{t}$ lie strictly within the unit circle—that is, when $\rho(\mathbf{A}_{t})<1$. In addition to the spectral radius $\rho$, we analyze the imaginary components of eigenvalues $\mathrm{Im}(\lambda_{i})$ to characterize oscillatory dynamics in the latent space. This helps reveal cycles or damped oscillations in control behavior, which may not be visible through magnitude alone. Specifically, local dynamics are characterized as follows:
  • $\rho(\mathbf{A}_{t})\leq 1$ with $\mathrm{Im}(\lambda_{i})=0$: non-oscillatory locally stable dynamics,
  • $\rho(\mathbf{A}_{t})\leq 1$ with $\mathrm{Im}(\lambda_{i})\neq 0$: locally stable dynamics with damped oscillations,
  • $\rho(\mathbf{A}_{t})>1$: locally unstable dynamics, regardless of the imaginary component.
While formal global stability of linear time-varying systems depends on the joint spectral radius (JSR) of the operator sequence $\{I+\mathbf{A}_{t}\}$ [24], computing the JSR is generally infeasible when $\mathbf{A}_{t}$ is state-dependent and varies continuously across time. SALSA-RL instead performs stability diagnostics locally at fixed state snapshots using the spectral radius of $\mathbf{A}_{t}$, offering a tractable and interpretable view of action evolution in specific regions of state space. Indeed, one may use this formulation to assess the spectral properties of the latent dynamics for various state-action combinations before deploying any pretrained agent.
Additionally, neighborhood states around the observed states $s_{\text{range}}$ are evaluated by computing the corresponding $\mathbf{A}_{t}$ and analyzing their eigenvalues and spectral radius. This reveals locally stable or unstable regions, capturing how state variations influence stability. Importantly, our goal is not to analyze the full closed-loop environment-agent system but to empirically characterize the behavior of the learned policy through its latent action dynamics. This local stability characterization serves as a proxy for inferring how consistent, structured, or potentially unstable the policy’s actions are over time, especially in high-stakes or safety-critical applications where local interpretability is crucial. This also provides insights into regions where the policy achieves stable control or displays unstable behavior. The algorithm is outlined in Appendix B.